DIFFERENTIAL GEOMETRY/PDE SEMINAR

Tuesday, July 22, 2003 Padelford C-36 1:30 pm (Note: Special Day and Time!)

Multi well problems and shape memory materials

Bernd Kirchheim (University of Oxford)

We discuss differential inclusions describing the zero energy state for certain shape memory alloys and discuss attempts and ideas to construct solutions for this inclusion.

One aim is to present recent progress in the question which global behaviour in 3 dimensions allows almost energy free mikrostructurs and the construction of particular simple solutions for a 2 dimensional model problem.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.